# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v4

@@ Abstract
Meta reinforcement learning (meta-RL) methods such as RL2 have emerged as promising approaches for learning data-efficient RL algorithms tailored to a given task distribution. However, they show poor asymptotic performance and struggle with out-of-distribution tasks because they rely on sequence models, such as recurrent neural networks or transformers, to process experiences rather than summarize them using general-purpose RL components such as value functions. In contrast, traditional RL algorithms are data-inefficient as they do not use domain knowledge, but they do converge to an optimal policy in the limit. We propose RL3, a principled hybrid approach that incorporates action-values, learned per task through traditional RL, in the inputs to meta-RL. We show that RL3 earns greater cumulative reward in the long term, compared to RL2, while maintaining data-efficiency in the short term, and generalizes better to out-of-distribution tasks. Experiments are conducted on both custom and benchmark discrete domains from the meta-RL literature that exhibit a range of short-term, long-term, and complex dependencies.
@@ 1 Introduction
Reinforcement learning (RL) has been shown to produce effective policies in a variety of applications including both virtual Mnih et al. (2015) and embodied Schulman et al. (2017); Haarnoja et al. (2018) systems. However, traditional RL algorithms have three major drawbacks: they often have difficulty generalizing beyond the exact task they were trained on, can be slow to converge, and require a large amount of data. These shortcomings are especially glaring in settings where the goal is to learn policies for a collection or distribution of problems that share some similarities, and for which traditional RL must start from scratch for each problem. For example, many robotic manipulation tasks require interacting with an array of objects with similar but not identical shapes, sizes, weights, materials, and appearances, such as mugs and cups. It is likely that effective manipulation strategies for these tasks will be similar, but they may also differ in ways that make it challenging to learn a single policy that is highly successful on all instances. Recently, meta reinforcement learning (meta-RL) has been proposed as an approach to mitigate these shortcomings by deriving RL algorithms (or meta-RL policies) that adapt efficiently to a distribution of tasks that share some common structure Duan et al. (2016); Wang et al. (2016).
Table 1: RL3 combines the strengths of meta-RL (e.g., RL2) and traditional RL. Like RL2, RL3 uses finite-context sequence models to represent data-efficient RL algorithms, optimized for tasks within a specified distribution. However, RL3 also includes a general-purpose RL routine that distills arbitrary amounts of data into optimal value-function estimates during adaptation. This improves long-term reasoning and out-of-distribution generalization.
@@
(Universal) (Improved)
While meta-RL systems represent a significant improvement over traditional RL in such settings, they still require large amounts of data during meta-training time, can have poor asymptotic performance during adaptation, and although they “learn to learn,” they often generalize poorly to tasks not represented in the meta-training distribution. This is partly because they rely on black-box sequence models like recurrent neural networks or transformers to process experience data. These models cannot handle arbitrary amounts of data effectively and lack integrated general-purpose RL components that could induce a broader generalization bias.
Hence, we propose RL3, a principled approach that embeds the strengths of traditional RL within meta-RL. Table 1 highlights our primary aims and the foremost insight informing our approach. The key idea in RL3 is an additional ‘object-level’ RL procedure executed within the meta-RL architecture that computes task-specific optimal Q-value estimates as supplementary inputs to the meta-learner, in conjunction with the sequence of states, actions and rewards. In principle, our approach allows the meta-learner to learn how to optimally fuse raw experience data with the summarizations provided by the Q-estimates. Ultimately, RL3 leverages the general-purpose nature of Q-value estimates, their ability to compress large amounts of experiences into useful summaries, their direct actionability, and their asymptotic optimality in order to enhance long-term performance and out-of-distribution (OOD) generalization. In this work, RL3 is implemented by injecting Q-value estimates into RL2 Duan et al. (2016).
The primary contribution of this paper is a confirmation of the hypothesis that injecting Q-estimates obtained via traditional object-level RL alongside the typical experience histories within a meta-RL agent leads to higher long-term returns and better OOD generalization, while maintaining short-term efficiency. We further demonstrate that our approach can also work with an abstract, or coarse, representation of the object-level MDP. Our conclusions are based on the results of experiments on the Bandits and MDPs domains used in previous work Duan et al. (2016); Mishra et al. (2018) as well as on a more challenging custom GridWorld domain that requires long-term reasoning. FInally, we examine the key insights that inform our approach and show theoretically that object-level Q-values are directly related to the optimal meta-value function.
@@ 2 Related Work
Although meta-RL is a fairly new topic of research, the general concept of meta-learning is decades old Vilalta and Drissi (2002), which, coupled with a significant number of design decisions for meta-RL systems, has created a large number of different proposals for how systems ought to best exploit the resources available within their deployment contexts Beck et al. (2023). At a high level, most meta-RL algorithms can be categorized as either parameterized policy gradient (PPG) models Finn et al. (2017); Li et al. (2017); Sung et al. (2017); Al-Shedivat et al. (2018); Gupta et al. (2018); Yoon et al. (2018); Stadie et al. (2018); Vuorio et al. (2019); Zintgraf et al. (2019); Raghu et al. (2019); Kaushik et al. (2020); Ghadirzadeh et al. (2021); Mandi et al. (2022) or black box models Duan et al. (2016); Heess et al. (2015); Wang et al. (2016); Foerster et al. (2018); Mishra et al. (2018); Humplik et al. (2019); Fakoor et al. (2020); Yan et al. (2020); Zintgraf et al. (2020); Liu et al. (2021); Emukpere et al. (2021); Beck et al. (2022). PPG approaches assume that the underlying learning process is best represented as a policy gradient, where the set of parameters that define the underlying algorithm ultimately form a differentiable set of meta-parameters that the meta-RL system may learn to adjust. The additional structure provided by this assumption, combined with the generality of policy gradient methods, means that typically PPG methods retain greater generalization capabilities on out-of-distribution tasks. However, due to their inherent data requirements, PPG methods are often slower to adapt and initially train.
@@
In this paper we focus on black box models, which represent the meta-learning function as a neural network, often a recurrent neural network (RNN) Duan et al. (2016); Heess et al. (2015); Wang et al. (2016); Humplik et al. (2019); Fakoor et al. (2020); Yan et al. (2020); Zintgraf et al. (2020); Liu et al. (2021) or a transformer Mishra et al. (2018); Wang et al. (2021); Melo (2022). There are also several hybrid approaches that combine PPG and black box methods, either during meta-training Ren et al. (2023) or fine-tuning Lan et al. (2019); Xiong et al. (2021). Using black box models simplifies the process of augmenting meta states with Q-estimates and allows us to retain relatively better data efficiency while relying on the Q-value injections for better long-term performance and generalization.
@@
Meta-RL systems may also leverage extra information available during training, such as task identification Humplik et al. (2019); Liu et al. (2021). Such ‘privileged information’ can of course lead to more performant systems, but is not universally available. As our hypothesis does not rely on the availability of such information, we expect our approach to be orthogonal to, and compatible with, such methods. Black box meta-RL systems that do not use privileged information still vary in several ways, including the choice between on-policy and off-policy learning and, in systems that use neural networks, the choice between transformers Vaswani et al. (2017) and RNNs Elman (1990); Hochreiter and Schmidhuber (1997); Cho et al. (2014).
@@
The most relevant methods to our work are end-to-end methods, which use a single function approximator to subsume both learner and meta-learner, such as RL2 Duan et al. (2016), L2L Wang et al. (2016), SNAIL Mishra et al. (2018), and E-RL2 Stadie et al. (2018), and methods that exploit the formal description of the meta-RL problem as a POMDP or a Bayes-adaptive MDP (BAMDP) Duff (2002). These methods attempt to learn policies conditioned on the BAMDP belief state while also approximating this belief state by, for example, variational inference (VariBAD) Zintgraf et al. (2020); Dorfman et al. (2020), or random network distillation on belief states (HyperX) Zintgraf et al. (2021). Or, they simply encode enough experience history to approximate POMDP beliefs (RL2) Duan et al. (2016); Wang et al. (2016).
@@ 3.1 Partially Observable MDPs
Markov decision processes (MDPs) underpin reinforcement learning, and in this paper we use the standard notation defining an MDP as a tuple $M=\langle S,A,T,R\rangle$, where $S$ is a set of states; $A$ is a set of actions; $T$ is the transition and $R$ is the reward function. A partially observable Markov decision process (POMDP) extends MDPs to settings with partially observable states. A POMDP is described as a tuple $\langle S,A,T,R,\Omega,O\rangle$, where $S,A,T,R$ are as in an MDP. $\Omega$ is the set of possible observations, and $O:S\times A\times\Omega\rightarrow[0,1]$ is an observation function representing the probability of receiving observation $\omega$ after performing action $a$ and transitioning to state $s^{\prime}$. POMDPs can alternatively be represented as continuous-state belief-MDPs where a belief state $b\in\Delta^{|S|}$ is a probability distribution over all states. In this representation, a policy $\pi$ is a mapping from belief states to actions, $\pi:\Delta^{|S|}\rightarrow A$.
@@ 3.2 Reinforcement Learning
Reinforcement learning (RL) agents learn an optimal policy given an MDP with unknown dynamics using only transition and reward feedback. This is often done by incrementally estimating the optimal action-value function $Q^{*}(s,a)$ Watkins and Dayan (1992), which satisfies the Bellman optimality equation $Q^{*}(s,a)=\mathbb{E}_{s^{\prime}}[R(s,a)+\gamma\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime})]$. In large or continuous state settings, it is popular to use deep neural networks to represent the action-value functions Mnih et al. (2015). We denote the vector representing the Q-estimates of all actions at state $s$ as $Q(s)$, and after $t$ feedback steps, as $Q^{t}(s)$. Q-learning is known to converge asymptotically Sutton and Barto (2018), provided each state-action pair is explored sufficiently. As a rough general statement, $||Q^{t}(s)-Q^{*}(s)||_{\infty}$ is proportional to $\approx\frac{1}{\sqrt{t}}$, with strong results on the convergence error available Szepesvári (1997); Kearns and Singh (1998); Even-Dar et al. (2003). The theoretical objective in RL is to optimize the value of the final policy, i.e. the cumulative reward per episode, disregarding the data cost incurred and the cumulative reward missed (or regret) during learning due to suboptimal exploration.
@@ 3.3 Meta Reinforcement Learning
Meta reinforcement learning seeks action selection strategies that minimize regret in MDPs drawn from a distribution of MDPs that share the same state and action spaces. Therefore, the objective in meta-RL is to maximize the cumulative reward over the entire interaction (or adaptation) period with an MDP, which may span multiple episodes, in order to optimize the exploration-exploitation tradeoff. Formally,
@@
Another way to conceptualize this problem is to recognize that the meta-RL problem may be written as a meta-level POMDP, where the hidden variable is the particular MDP (or task) at hand, $M_{i}$, which varies across meta-episodes. This framing, known as Bayesian RL Ghavamzadeh et al. (2015), leverages the fact that augmenting the task-specific state $s$ with belief over tasks $b(i)$ results in a Markovian meta-state $[s,b]$ for optimal action selection, a model known as the Bayes Adaptive MDP (or BAMDP) Duff (2002). That is, this belief state captures all requisite information for the purpose of acting. We will revisit this concept to develop intuition on the role of object-level Q-value estimates in the meta-RL value function.
@@ 4 RL3
To address the limitations of black box meta-RL methods, we propose RL3, a principled approach that leverages (1) the inherent generality of action-value estimates, (2) their ability to compress experience histories into useful summaries, (3) their direct actionability & asymptotic optimality, (4) their ability to inform task-identification, and (5) their relation to the optimal meta-value function, in order to enhance out-of-distribution (OOD) generalization and performance over extended adaptation periods. The central, novel mechanism in RL3 is an additional ‘object-level’ RL procedure executed within the meta-RL architecture, shown in Fig. 1, that computes task-specific optimal Q-value estimates $Q^{t}_{i}(s_{t})$ and state-action counts as supplementary inputs to the meta-RL policy in conjunction with the sequence of states, actions and rewards $(s_{0},a_{0},r_{0},...,s_{t})$. The Q-estimates are computed off-policy, and may involve model estimation and planning for greater data efficiency. The estimates and the counts are reset at the beginning of each meta-episode as a new task $M_{i}$ is sampled. In all subsequent text, Q-value estimates used as input entail the inclusion of state-action counts as well. We now present a series of key insights informing our approach.
@@
First, estimating action-values is a key component in many universal RL algorithms, and asymptotically, they fully inform optimal behavior irrespective of domain. Strategies for optimal exploration-exploitation trade-off are domain-dependent and rely on historical data, yet many exploration approaches use estimated Q-values and some notion of counts alone, such as epsilon-greedy, Boltzmann exploration, upper confidence bounds (UCB/UCT) Auer (2002); Kocsis and Szepesvári (2006), count-based exploration Tang et al. (2017), curiosity based exploration Burda et al. (2019) and maximum-entropy RL Haarnoja et al. (2018). This creates a strong empirical case that using Q-value estimates and state-action counts for efficient exploration has inherent generality.
Figure 2: Sub-figure (a) shows a meta-episode in a shortest-path environment where the goal position (green circles) and the obstacles (black regions) may vary across tasks. In this meta-episode, after the meta-RL agent narrows its belief about the goal position of this task (dark-green circle) having followed a principled exploration strategy ($\tau_{0}$), it explores potential shorter paths in subsequent episodes ($\tau_{1},\tau_{2},\tau_{3},\tau_{4}$). Throughout this process, the estimated value-function $\hat{Q}^{*}$ implicitly “remembers” the goal position and previous paths traversed in a finite-size representation, and updates the shortest path calculation (highlighted in bold) using Bellman backups when paths intersect. Sub-figures (b) and (c) illustrate the many-to-one mapping of object- and meta-level data streams to Q-estimates, and thus their utility as compression and summarization mechanisms for meta-learning.
@@
Second, Q-estimates summarize experience histories of arbitrary length and order in one constant-size vector. This mapping is many-to-one, and any permutation of transitions ($\langle s,a,r,s^{\prime}\rangle$ tuples) or episodes in a history of experiences yield the same Q-estimates. Although this compression is lossy, it still “remembers” important aspects of the experienced episodes, such as high-return actions and goal positions (see Fig. 2) since Q-estimates persist across episodes. This simplifies the mapping the meta agent needs to learn as Q-estimates represent a smaller and more salient set of inputs compared to all possible histories with the same implication.
@@
Fourth, Q-estimates are excellent task discriminators and serve as another line of evidence vis-à-vis maintaining belief over tasks. In a simple domain like Bernoulli multi-armed bandits Duan et al. (2016), Q-estimates and action-counts combined are sufficient for Bayes-optimal behavior even without providing raw experience data – a result surprisingly unstated in the literature to the best of our knowledge (see Appendix A.1). However, Q-estimates and action-counts may not always be sufficient for Bayes-optimal beliefs. For example, in Gaussian multi-armed bandits, the sufficient statistics include the variance in rewards for each action (see Appendix A.2). In more complex domains, it is hard to prove the sufficiency of Q-estimates regarding task discrimination. However, via empirical analysis in Appendix D, we argue that i) it is highly improbable for two tasks to have similar Q functions and ii) Q-estimates tend to become accurate task predictors in just a few steps. This implies that the meta-agent may use this finite summary for task inference rather than relying completely on arbitrarily long histories, potentially contributing to enhanced performance over long adaptation periods.
@@
It can be theoretically argued that since the meta agent is a BAMDP policy, it is meta-trained to select greedy actions w.r.t. the BAMDP meta-value function and thus should not require constructing a task-specific plan internally. However, the optimality of the meta action-value function depends on implicitly (or explicitly in some approaches, such as Humplik et al. (2019); Zintgraf et al. (2020); Dorfman et al. (2020); Zintgraf et al. (2021)) maintaining a Bayes-optimal belief over tasks in the transformer/RNN architecture. This may be challenging if the task distribution is too broad and the function approximator is not powerful enough to integrate experience histories into Bayes-optimal beliefs, or altogether impossible if there is a distribution shift at meta-test time. This latter condition is common in practice and is a frequent target use case for meta-RL systems. Incorporating task-specific Q-estimates gives the agent a simple alternative (even if not Bayes-optimal) line of reasoning to translate experiences into actions. Incorporating Q-estimates thus reduces susceptibility to distribution shifts since the arguments presented in this section are domain independent.
@@
Finally, Q-estimates often converge far more quickly than the theoretical rate of $\frac{1}{\sqrt{t}}$, allowing them to be useful in the short and medium term, since i) most real-world domains contain significant determinism, ii) it is not necessary to estimate Q-values for states unreachable by the optimal policy, and iii) optimal meta-RL policies may represent active exploration strategies in which Q-estimates converge faster, or evolve in a manner leading to quicker task identification. This is intuitively apparent in shortest-path problems, as illustrated in Fig. 2(a). In a deep neural network, it is difficult to know exactly how Q-estimates will combine with state-action-reward histories when approximating the meta-value function. However, as we show below, we can rather straightforwardly write an equation for the meta-value function in terms of these constituent streams of information, which may explain why this function is seemingly relatively easy to learn compared to predicting meta-values from histories alone.

### v5

@@ Abstract
Meta reinforcement learning (meta-RL) methods such as RL2 have emerged as promising approaches for learning data-efficient RL algorithms tailored to a given task distribution. However, they show poor asymptotic performance and struggle with out-of-distribution tasks because they rely on sequence models, such as recurrent neural networks or transformers, to process experiences rather than summarize them using general-purpose RL components such as value functions. In contrast, traditional RL algorithms are data-inefficient as they do not use domain knowledge, but do converge to an optimal policy in the limit. We propose RL3, a principled hybrid approach that incorporates action-values, learned per task via traditional RL, in the inputs to meta-RL. We show that RL3 earns greater cumulative reward in the long term compared to RL2 while drastically reducing meta-training time and generalizes better to out-of-distribution tasks. Experiments are conducted on both custom and benchmark discrete domains from the meta-RL literature that exhibit a range of short-term, long-term, and complex dependencies.
@@ 1 Introduction
Reinforcement learning (RL) has been shown to produce effective policies in a variety of applications including both virtual [1] and embodied [2, 3] systems. However, traditional RL algorithms have three major drawbacks: they can be slow to converge, require a large amount of data, and often have difficulty generalizing to out-of-distribution (OOD) tasks not practiced during training. These shortcomings are especially glaring in settings where the goal is to learn policies for a collection or distribution of problems that share some similarities, and for which traditional RL must start from scratch for each problem. For example, many robotic manipulation tasks require interacting with an array of objects with similar but not identical shapes, sizes, weights, materials, and appearances, such as mugs and cups. It is likely that effective manipulation strategies for these tasks will be similar, but they may also differ in ways that make it challenging to learn a single policy that is highly successful on all instances. Recently, meta reinforcement learning (meta-RL) has been proposed as an approach to mitigate these shortcomings by deriving RL algorithms (or meta-RL policies) that adapt efficiently to a distribution of tasks that share some common structure [4, 5].
While meta-RL systems represent a significant improvement over traditional RL in such settings, they still require large amounts of data during meta-training time, can have poor asymptotic performance during adaptation, and although they “learn to learn," they often generalize poorly to tasks not represented in the meta-training distribution. This is partly because they rely on black-box sequence models like recurrent neural networks or transformers to process experience data. These models cannot handle arbitrary amounts of data effectively and lack integrated general-purpose RL components that could induce a broader generalization bias.
Table 1: RL3 combines the strengths of RL2 and traditional RL. Like RL2, RL3 uses finite-context sequence models to represent data-efficient RL algorithms, optimized for tasks within a specified distribution. However, RL3 also includes a general-purpose RL routine that distills arbitrary amounts of data into optimal value-function estimates during adaptation. This improves long-term reasoning and OOD generalization.
@@
(General Purpose) (Improved)
Hence, we propose RL3, an approach that embeds the strengths of traditional RL within meta-RL. Table 1 highlights our primary aims and the foremost insight informing our approach. The key idea in RL3 is an additional ‘object-level’ RL procedure executed within the meta-RL architecture that computes task-specific optimal Q-value estimates as supplementary inputs to the meta-learner, in conjunction with sequences of states, actions and rewards. In principle, our approach allows the meta-learner to learn how to optimally fuse raw experience data with summarizations provided by the Q-estimates. Ultimately, RL3 leverages Q-estimates’ generality, ability to compress large amounts of experiences into useful summaries, direct actionability, and asymptotic optimality to enhance long-term performance and OOD generalization and drastically reduce meta-training time.
While Q-value estimates can be injected into any other meta-RL algorithm, for clarity of exposition, we implement RL3 by injecting Q-value estimates into one of the most popular and easily understood meta-RL algorithm, RL2 [4] (hence, the name RL3). However, it should be noted that our baseline implementation of RL2 includes significant enhancements like using transformers instead of LSTMs to improve long-context reasoning, in addition to incorporating numerous recommendations from Ni et al. [6] that have been shown to make recurrent model-free RL algorithms like RL2 competitive with state-of-the-art meta-RL baselines like VariBAD [7].
The primary contribution of this paper is a proof-of-concept that injecting Q-estimates obtained via traditional object-level RL alongside the typical experience histories within a meta-RL agent leads to higher long-term returns and better OOD generalization, while maintaining short-term efficiency. We further demonstrate that our approach can also work with an abstract, or coarse, representation of the object-level MDP. We experiment with discrete domains that both reflect the challenges faced by meta-RL and simultaneously allow transparent analysis of the results. Finally, we examine the key insights that inform our approach and show theoretically that object-level Q-values are directly related to the optimal meta-value function.
@@ 2 Related Work
Although meta-RL is a fairly new topic of research, the general concept of meta-learning is decades old [8], which, coupled with a significant number of design decisions for meta-RL systems, has created a large number of different proposals for how systems ought to best exploit the resources available within their deployment contexts [9]. At a high level, most meta-RL algorithms can be categorized as either parameterized policy gradient (PPG) models [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] or black box models [4, 23, 5, 24, 25, 26, 27, 28, 7, 29, 30, 31]. PPG approaches assume that the underlying learning process is best represented as a policy gradient, where the set of parameters that define the underlying algorithm ultimately form a differentiable set of meta-parameters that the meta-RL system may learn to adjust. The additional structure provided by this assumption, combined with the generality of policy gradient methods, means that typically PPG methods retain greater generalization capabilities on out-of-distribution tasks. However, due to their inherent data requirements, PPG methods are often slower to adapt and initially train.
@@
In this paper we focus on black box models, which represent the meta-learning function as a neural network, often a recurrent neural network (RNN) [4, 23, 5, 26, 27, 28, 7, 29] or a transformer [25, 32, 33]. There are also several hybrid approaches that combine PPG and black box methods, either during meta-training [34] or fine-tuning [35, 36]. Using black box models simplifies the process of augmenting meta states with Q-estimates and allows us to retain relatively better data efficiency while relying on the Q-value injections for better long-term performance and generalization.
@@
Meta-RL systems may also leverage extra information available during training, such as task identification [26, 29]. Such ‘privileged information’ can of course lead to more performant systems, but is not universally available. As our hypothesis does not rely on the availability of such information, we expect our approach to be orthogonal to, and compatible with, such methods. Black box meta-RL systems that do not use privileged information still vary in several ways, including the choice between on-policy and off-policy learning and, in systems that use neural networks, the choice between transformers [37] and RNNs [38, 39, 40].
@@
The most relevant methods to our work are end-to-end methods, which use a single function approximator to subsume both learner and meta-learner, such as RL2 [4], L2L [5], SNAIL [25], and E-RL2 [16], and methods that exploit the formal description of the meta-RL problem as a POMDP or a Bayes-adaptive MDP (BAMDP) [41]. These methods attempt to learn policies conditioned on the BAMDP belief state while also approximating this belief state by, for example, variational inference (VariBAD) [7, 42], or random network distillation on belief states (HyperX) [43]. Or, they simply encode enough experience history to approximate POMDP beliefs (RL2) [4, 5].
@@ 3.1 Partially Observable MDPs
We use the standard notation defining a Markov decision process (MDP) as a tuple $M=\langle S,A,T,R\rangle$, where $S$ is a set of states; $A$ is a set of actions; $T$ is the transition and $R$ is the reward function. A partially observable Markov decision process (POMDP) extends MDPs to settings with partially observable states. A POMDP is described as a tuple $\langle S,A,T,R,\Omega,O\rangle$, where $S,A,T,R$ are as in an MDP. $\Omega$ is the set of possible observations, and $O:S\times A\times\Omega\rightarrow[0,1]$ is an observation function representing the probability of receiving observation $\omega$ after performing action $a$ and transitioning to state $s^{\prime}$. POMDPs can alternatively be represented as continuous-state belief-MDPs where a belief state $b\in\Delta^{|S|}$ is a probability distribution over all states. In this representation, a policy $\pi$ is a mapping from belief states to actions, $\pi:\Delta^{|S|}\rightarrow A$.
@@ 3.2 Reinforcement Learning
Reinforcement learning (RL) agents learn an optimal policy given an MDP with unknown dynamics using only transition and reward feedback. This is often done by incrementally estimating the optimal action-value function $Q^{*}(s,a)$ [44], which satisfies the Bellman optimality equation $Q^{*}(s,a)=\mathbb{E}_{s^{\prime}}[R(s,a)+\gamma\max_{a^{\prime}\in A}Q^{*}(s^{\prime},a^{\prime})]$. In large or continuous state settings, it is popular to use deep neural networks to represent the action-value functions [1]. We denote the vector representing the Q-estimates of all actions at state $s$ as $Q(s)$, and after $t$ feedback steps, as $Q^{t}(s)$. Q-learning is known to converge asymptotically [45], provided each state-action pair is explored sufficiently. As a rough general statement, $||Q^{t}(s)-Q^{*}(s)||_{\infty}$ is proportional to $\approx\frac{1}{\sqrt{t}}$, with strong results on the convergence error available [46, 47, 48]. The theoretical objective in RL is to optimize the value of the final policy i.e., the cumulative reward per episode, disregarding the data cost incurred and the cumulative reward missed (or regret) during learning due to suboptimal exploration.
@@ 3.3 Meta Reinforcement Learning
Meta-RL seeks action selection strategies that minimize regret in MDPs drawn from a distribution of MDPs that share the same state and action spaces. Therefore, the objective in meta-RL is to maximize the cumulative reward over the entire interaction (or adaptation) period with an MDP, which may span multiple episodes, in order to optimize the exploration-exploitation tradeoff. Formally,
@@
Another way to conceptualize this problem is to recognize that the meta-RL problem may be written as a meta-level POMDP, where the hidden variable is the particular MDP (or task) at hand, $M_{i}$, which varies across meta-episodes. This framing, known as Bayesian RL [49], leverages the fact that augmenting the task-specific state $s$ with belief over tasks $b(i)$ results in a Markovian meta-state $[s,b]$ for optimal action selection, a model known as the Bayes Adaptive MDP (or BAMDP) [41]. That is, this belief state captures all requisite information for the purpose of acting. We will revisit this concept to develop intuition on the role of object-level Q-value estimates in the meta-RL value function.
@@ 4 RL3
To address the limitations of black box meta-RL methods, we propose RL3, a principled approach that leverages (1) the inherent generality of action-value estimates, (2) their ability to compress experience histories into useful summaries, (3) their direct actionability & asymptotic optimality, (4) their ability to inform task-identification, and (5) their relation to the optimal meta-value function, in order to enhance out-of-distribution (OOD) generalization and performance over extended adaptation periods. The central, novel mechanism in RL3 is an additional ‘object-level’ RL procedure executed within the meta-RL architecture, shown in Fig. 1, that computes task-specific optimal Q-value estimates $Q^{t}_{i}(s_{t})$ and state-action counts as supplementary inputs to the meta-RL policy in conjunction with the sequence of states, actions and rewards $(s_{0},a_{0},r_{0},...,s_{t})$. The Q-estimates are computed off-policy, and may involve model estimation and planning, for greater data efficiency. The estimates and the counts are reset at the beginning of each meta-episode as a new task $M_{i}$ is sampled. In all subsequent text, Q-value estimates used as input entail the inclusion of state-action counts as well. We now present a series of key insights informing our approach.
@@
First, estimating action-values is a key component in many universal RL algorithms, and asymptotically, they fully inform optimal behavior irrespective of domain. Strategies for optimal exploration-exploitation trade-off are domain-dependent and rely on historical data, yet many exploration approaches use estimated Q-values and some notion of counts alone, such as epsilon-greedy, Boltzmann exploration, upper confidence bounds (UCB/UCT) [50, 51], count-based exploration [52], curiosity based exploration [53] and maximum-entropy RL [3]. This creates a strong empirical case that using Q-value estimates and state-action counts for efficient exploration has inherent generality.
@@
Second, Q-estimates summarize experience histories of arbitrary length and order in one constant-size vector. This mapping is many-to-one, and any permutation of transitions ($\langle s,a,r,s^{\prime}\rangle$ tuples) or episodes in a history of experiences yield the same Q-estimates. Although this compression is lossy, it still “remembers" important aspects of the experienced episodes, such as high-return actions and goal positions (see Fig. 2) since Q-estimates persist across episodes. This simplifies the mapping the meta agent needs to learn as Q-estimates represent a smaller and more salient set of inputs compared to all possible histories with the same implication.
@@
Figure 2: Sub-figure (a) shows a meta-episode in a shortest-path environment where the goal position (green circles) and the obstacles (black regions) may vary across tasks. In this meta-episode, after the meta-RL agent narrows its belief about the goal position of this task (dark-green circle) having followed a principled exploration strategy ($\tau_{0}$), it explores potential shorter paths in subsequent episodes ($\tau_{1},\tau_{2},\tau_{3},\tau_{4}$). Throughout this process, the estimated value-function $\hat{Q}^{*}$ implicitly “remembers" the goal position and previous paths traversed in a finite-size representation, and updates the shortest path calculation (highlighted in bold) using Bellman backups when paths intersect. Sub-figures (b) and (c) illustrate the many-to-one mapping of object- and meta-level data streams to Q-estimates, and thus their utility as compression and summarization mechanisms for meta-learning.
@@
Fourth, Q-estimates are excellent task discriminators and serve as another line of evidence vis-à-vis maintaining belief over tasks. In a simple domain like Bernoulli multi-armed bandits [4], Q-estimates and action-counts combined are sufficient for Bayes-optimal behavior even without providing raw experience data – a result surprisingly unstated in the literature to the best of our knowledge (see Appendix A.1). However, Q-estimates and action-counts may not always be sufficient for Bayes-optimal beliefs [^1] For example, in Gaussian multi-armed bandits, the sufficient statistics include the variance in rewards for each action (see Appendix A.2).. In more complex domains, it is hard to prove the sufficiency of Q-estimates regarding task discrimination. However, via empirical analysis in Appendix D, we argue that i) it is highly improbable for two tasks to have similar Q functions and ii) Q-estimates tend to become accurate task predictors in just a few steps. This implies that the meta-agent may use this finite summary for task inference rather than relying completely on arbitrarily long histories, potentially contributing to enhanced performance over long adaptation periods.
@@
It can be theoretically argued that since the meta agent is a BAMDP policy, it is meta-trained to select greedy actions w.r.t. the BAMDP meta-value function and thus should not require constructing a task-specific plan internally. However, the optimality of the meta action-value function depends on implicitly (or explicitly in some approaches [26, 7, 42, 43]) maintaining a Bayes-optimal belief over tasks in the transformer/RNN architecture. This may be challenging if the task distribution is too broad and the function approximator is not powerful enough to integrate experience histories into Bayes-optimal beliefs, or altogether impossible if there is a distribution shift at meta-test time. This latter condition is common in practice and is a frequent target use case for meta-RL systems. Incorporating task-specific Q-estimates gives the agent a simple alternative (even if not Bayes-optimal) line of reasoning to translate experiences into actions. Incorporating Q-estimates thus reduces susceptibility to distribution shifts since the arguments presented in this section are domain independent.
@@
Finally, Q-estimates often converge far more quickly than the theoretical rate of $\frac{1}{\sqrt{t}}$, allowing them to be useful in the short and medium term, since i) most real-world domains contain significant determinism, ii) it is not necessary to estimate Q-values for states unreachable by the optimal policy, and iii) optimal meta-RL policies may represent active exploration strategies in which Q-estimates converge faster, or evolve in a manner leading to quicker task identification. This is intuitively apparent in shortest-path problems, as illustrated in Fig. 2(a). In a deep neural network, it is difficult to know exactly how Q-estimates will combine with state-action-reward histories when approximating the meta-value function. However, as we show below, we can write an equation for the meta-value function in terms of these constituent streams of information, which may explain why this function is seemingly relatively easy to learn compared to predicting meta-values from histories alone.